Solve the nonlinear inequality. Express the solution using interval notation.
x/2 ≥ 5/(x + 1)+ 4

Graph the solution
Answer #1

Given inequality is:

\frac{x}{2}\geq \frac{5}{x+1}+4

\frac{x}{2}\geq \frac{5+4(x+1)}{x+1}

\frac{x}{2}\geq \frac{5+4x+4}{x+1}

\frac{x}{2}\geq \frac{9+4x}{x+1}

x(x+1)\geq 2(9+4x)

x^2+x\geq 18 +8x

x^2+x-8x-18\geq 0

x^2-7x-18\geq 0

x^2-9x+2x-18\geq 0

x(x-9)+2(x-9)\geq 0

(x-9)(x+2)\geq 0

So solutions of given inequalities are:

x\in (-\infty ,-2]\cup [9,\infty )